A physician orders lorazepam (Ativan) 0.05 mg/kg I.M. two hours prior to a procedure

Question 4 

A physician orders lorazepam (Ativan) 0.05 mg/kg I.M. two hours prior to a procedure. The lorazepam is supplied in a vial containing 2 mg/mL. The patient weighs 50 kg. How many mL should the nurse give? 


Question 5 

A nurse receives the following order from the physician: Give 5,000 units heparin sodium subcutaneously STAT. The heparin is available in a vial containing 10,000 units/mL. How many mL should the nurse give? 


Question 6 

The next order the nurse receives reads: After the initial heparin bolus, give 9.000 units of heparin sodium subcutaneously TID. The heparin in available in a vial containing 5,000  unit/mL. How many mL would the nurse give in one dose?
Answer #1

Q.4. ordered lorazepam= 0.05 mg/kg IM

supplied 2mg/ml

patient wt=50kg

as per wt dose will be= 0.05 X50= 2.5 mg

the required dose in ml?

2mg in 1ml

1mg in 1/2 ml

2.5 mg will be supplied from=2.5/2=1.25 ml answer

Q.5. order 5000 units of Heparin SC

available in 10000 units/ml

required heparin in ml=?

10000 units in 1ml

then 5000 will be in 0.5 ml answer

Q.6. required dose of heparin= 9000 units

available=5000units/ml

required dose in ml=?

5000 units in 1 ml

1 unit in =1/5000 ml

then 9000 units will be supplied from=(1/5000) X9000

=1.8 ml answer